OpenStudy (anonymous):

If y=5x, then the average ( arithmetic mean) of x and y, in term of x, is equal to A) x B) 2x C) 3x D) 4x E) 5x

OpenStudy (amistre64):

well; if x=2 then y=10 2+10 ----- = 6 2 3x in that case if x=4 y =20 3x = 12 24/2 = 12 id say 3x

OpenStudy (anonymous):

i would too

OpenStudy (anonymous):

thank you guys :)

OpenStudy (anonymous):

\[\frac{x+5x}{2}=3x\]

OpenStudy (anonymous):

so all you have to do is sub in any number for y

OpenStudy (amistre64):

satelitte always does it the easy way lol

OpenStudy (amistre64):

i lack the brains to go easy; i tend to brute math it all

OpenStudy (anonymous):

but it easy under stand lol xD